Order: Procainamide 5 mg/min IV drip

Available: Procainamide 2 g in 500 mL D5W
The IV flow rate for this dosage is how many mLs per hour?
Answer #1

Answer ;

The ordered amount of the drug is 5 mg/min, it can be written as 300 mg / hr
The drug which is available in the hand is 2 gm in 500 ml D5W, it means 2000 mg in 5OO ml,

So 1 ml contain = 2000 / 500
= 4 mg

We need to provide 300 mg/ hr, hence it is = 300 / 4
= 75

Therefore the dosage is 75 ml / hour.
Answer #2
5mgx60min=300/2000mgx 250ml=37.5 round to 38ml/hr 2g=2000mg
source: Dosage calculation nursing book
answered by: Anonymous